The students of same age group from two different schools were compared for variability in their mathematical skill. A random sample of 25 pupils from one school had a variance of 16 marks while a random sample of 22 pupils from the other school had a variance of 8 marks. Examine if the difference in variability is significant. [F at (24,21) at 5% level of significance is 2.05]


1
Expert's answer
2021-12-10T08:52:04-0500

The following null and alternative hypotheses need to be tested:

"H_0:\sigma_1^2\leq\sigma_2^2"

"H_1:\sigma_1^2>\sigma_2^2"

This corresponds to a right-tailed test, for which a F-test for two population variances needs to be used.

Numerator degrees of freedom: "d_1=25-1=24."

Denomirator degrees of freedom: "d_2=22-1=21."

Based on the information provided, the significance level is "\alpha = 0.05," and the the rejection region for this right-tailed test test is

"R = \{F: F > 2.054\}."

The F-statistic is computed as follows:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{16}{8}=2"

Since from the sample information we get that "F = 2 \le F_c = 2.054,"

it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population variance "\sigma_1^2" is greater than the population variance "\sigma_2^2," at the "\alpha = 0.05" significance level.